# TREASUREHUNT: Transients and Variability Discovered with HST in the JWST North Ecliptic Pole Time Domain Field

### v2

@@ 6 Conclusion
This study highlights the JWST NEP TDF as an exceptional site for investigating time-variable astronomical phenomena, including SNe and (weak) variable AGN. We identified 12 transients and 100 variable sources in ACS/WFC F606W and F435W images from the HST TREASUREHUNT program. We provide positional information, dates of observation, intervening time interval, and magnitudes for each of the 12 transients, which range in brightness from 24.8 to 27.5 mag. Their areal density is 0.07 transients per arcmin (245 per deg) per epoch. We argue that the vast majority of these transients are SNe. Three transients (T7, T10, and T11) were detected in X-rays, of which two (T7 and T10) appear isolated and point-like and are likely newly identified quasars. Transient T11 has no visible host galaxy, and its nature remains uncertain. We suspect it to be a faint quasar if the X-ray detection is indeed associated, or otherwise either part of a very faint dwarf galaxy at 6, or a transient in a more massive host at 6. One transient (T4) has a spectroscopic redshift = 0.615 from MMT/Binospec.
Our variability search revealed that 0.42% of the general 6 field galaxy population exhibits variability at 3 significance to depths of 29.5 mag in F606W and 28.6 mag in F435W, with an areal density of 1.25 variables per arcmin (4500 per deg). We carefully identified variable candidates, using the measured distribution of magnitude differences for objects observed more than once to calibrate our photometric uncertainties, where the scaled photometric uncertainty is , as defined in Equation 1. Sources are flagged as variable if they varied by more than in F435W or F606W, or varied by more than 2 in both filters in the same sense. This revealed 190 unique variable candidates, of which we estimated 80 to be false positives (using Gaussian statistics). Most of the variable candidates are coincident with the cores of galaxies, indicating potential AGN variability, while a smaller fraction appears associated with faint SNe. We also estimate photometric redshifts for our sample, from which we estimate masses, ages, dust extinction, and sSFRs. We briefly explore ways to estimate SMBH mass and AGN timescales using the time interval between overlapping visits and the observed rates of variability.
@@ Appendix A Assessment of the Potential Impact of Cosmic Rays on Detected Variability
A good recent summary of over 1.2 billion CRs analyzed during the 30 year operational lifetime of these four cameras inside HST is given by Miles et al. (e.g., 2021). Their Table 6 shows that both ACS, WFPC2, and WFC3 have a remarkably similar density of CR-hits of 1.10.1 hits/sec/cm of detector area. Miles et al. (2021) provide a better physically motivated Landau distribution of the CR-energies for all HST cameras, which resembles the earlier Weibull distribution of McMaster & Biretta (2010). These CR energy distributions are thus representative of the typical CR-distribution and frequency in typical HST images, as long as one stays away from transitions through the South Atlantic Anomaly (SAA), where the CR-flux can be temporarily much higher. Our NEP TDF images always avoid the SAA transitions, and therefore, we will assume that the CR-flux is as given in these respective Instrument Handbooks and Miles et al. (2021).
Following the IHB’s, at minimum four independent exposures of 1200 sec are needed to minimize multiple CR impacts on the same object, and we use a much more conservative N=8 exposures per pointing in each filter. Windhorst et al. (1994) investigate from 12-exposure pointings how much the CR-rejection improves if fewer than N=12 exposures are stacked, and find that for N8 the sky-mean and sky-sigma converge to the best achievable values (from the N=12 maximum exposure case) if one clips at the level of 3–3.5, where is the local sky rms noise-level. So starting with CR-clipping on 8 exposures for every pointing should be robust.

### v3

@@ 6 Conclusion
This study highlights the JWST NEP TDF as an exceptional site for investigating time-variable astronomical phenomena, including SNe and (weak) variable AGN. We identified 12 transients and 100 variable sources in ACS/WFC F606W and F435W images from the HST TREASUREHUNT program. We provide positional information, dates of observation, intervening time interval, and magnitudes for each of the 12 transients, which range in brightness from 24.8 to 27.5 mag. Their areal density is 0.07 transients per arcmin2 (245 per deg2) per epoch. We argue that the vast majority of these transients are SNe. Three transients (T7, T10, and T11) were detected in X-rays, of which two (T7 and T10) appear isolated and point-like and are likely newly identified quasars. Transient T11 has no visible host galaxy, and its nature remains uncertain. We suspect it to be a faint quasar if the X-ray detection is indeed associated, or otherwise either part of a very faint dwarf galaxy at 6, or a transient in a more massive host at 6. One transient (T4) has a spectroscopic redshift = 0.615 from MMT/Binospec.
Our variability search revealed that 0.42% of the general 6 field galaxy population exhibits variability at 3 significance to depths of 29.5 mag in F606W and 28.6 mag in F435W, with an areal density of 1.25 variables per arcmin2 (4500 per deg2). We carefully identified variable candidates, using the measured distribution of magnitude differences for objects observed more than once to calibrate our photometric uncertainties, where the scaled photometric uncertainty is , as defined in Equation 1. Sources are flagged as variable if they varied by more than in F435W or F606W, or varied by more than 2 in both filters in the same sense. This revealed 190 unique variable candidates, of which we estimated 80 to be false positives (using Gaussian statistics). Most of the variable candidates are coincident with the cores of galaxies, indicating potential AGN variability, while a smaller fraction appears associated with faint SNe. We also estimate photometric redshifts for our sample, from which we estimate masses, ages, dust extinction, and sSFRs. We briefly explore ways to estimate SMBH mass and AGN timescales using the time interval between overlapping visits and the observed rates of variability.
@@ Appendix A Assessment of the Potential Impact of Cosmic Rays on Detected Variability
A good recent summary of over 1.2 billion CRs analyzed during the 30 year operational lifetime of these four cameras inside HST is given by, e.g., Miles et al. (2021). Their Table 6 shows that both ACS, WFPC2, and WFC3 have a remarkably similar density of CR-hits of 1.10.1 hits/sec/cm2 of detector area. Miles et al. (2021) provide a better physically motivated Landau distribution of the CR-energies for all HST cameras, which resembles the earlier Weibull distribution of McMaster & Biretta (2010). These CR energy distributions are thus representative of the typical CR-distribution and frequency in typical HST images, as long as one stays away from transitions through the South Atlantic Anomaly (SAA), where the CR-flux can be temporarily much higher. Our NEP TDF images always avoid the SAA transitions, and therefore, we will assume that the CR-flux is as given in these respective Instrument Handbooks and Miles et al. (2021).
Following the Instrument Handbooks, at minimum four independent exposures of 1200 sec are needed to minimize multiple CR impacts on the same object, and we use a much more conservative N=8 exposures per pointing in each filter. Windhorst et al. (1994) investigate from 12-exposure pointings how much the CR-rejection improves if fewer than N=12 exposures are stacked, and find that for N8 the sky-mean and sky-sigma converge to the best achievable values (from the N=12 maximum exposure case) if one clips at the level of 3–3.5, where is the local sky rms noise-level. So starting with CR-clipping on 8 exposures for every pointing should be robust.